at 20 degrees celsius, a steel rod and an aluminum rod are both 1.25 m long. (a) if both rods are heated to 100 degrees, which will be longer and by how much?


1
Expert's answer
2021-02-10T10:07:22-0500

The elongation of a metal rod under heating is given as follows:


"\Delta L = \alpha L\Delta T"

where "\alpha" is the coefficient of thermal expansion of metal, "L" is the initial length, and "\Delta T" is the change in temperature. In our case: "L = 1.25m", "\Delta T = 100\degree C - 20\degree C = 80\degree C = 80K", the expansivities are:


"\alpha_{Al} = 23.1\times 10^{-6} K^{-1}\\
\alpha_{Cu} = 17\times 10^{-6} K^{-1}"

Thus, the elongation of aluminium is:


"\Delta L_{Al} = 23.1\times 10^{-6}\cdot 1.25\cdot 80 = 2.31mm"

The elongation of cooper is:


"\Delta L_{Cu} = 17\times 10^{-6}\cdot 1.25\cdot 80 =1.70mm"

Thus, the aluminum rod will be longer by


"\Delta L_{Al} - \Delta L_{Cu} = 2.31mm-1.70mm = 0.61mm"

Answer. The aluminum rod will be longer by 0.61 mm.